# A Unified Voxel Diffusion Module for Point Cloud 3D Object Detection

### v1

@@ Introduction
Ablation on Fine-grained Feature Aggregation. To further evaluate the fine-grained feature aggregation capability of VDM under more detailed voxel partitioning, we conducted experiments by refining the voxel resolution. Specifically, we reduced the voxel size along the x and y axes to one-fourth of the original size, while keeping the resolution along the z-axis unchanged, with voxel diffusion still enabled.
@@
As shown in Table 2, under this fine-grained aggregation setting, the model achieves a 0.5 APH (L2) improvement in detecting the Pedestrian category, indicating enhanced localization and recognition of small-scale objects. Furthermore, as reported in Table 7, on the ONCE dataset, the VDM-OD model exhibits inferior performance across all three categories when only voxel diffusion is applied. With the integration of fine-grained feature aggregation, the model achieves a 1.5 mAP gain, demonstrating the effectiveness of finer voxel partitioning in enhancing detection performance. When trained with 20% of the Waymo validation set, the VDM model also achieves higher detection accuracy for the Cyclist category.
Conclusion
We propose the Voxel Diffusion Module (VDM), which consists of a sparse 3D convolutional network, sparse residual blocks, and submanifold sparse convolution (SubM3D). This module diffuses voxelized point clouds via sparse 3D convolutions to increase the density of foreground voxels for more accurate object localization, while simultaneously performing spatial pre-aggregation of voxel features to provide spatial context for the subsequent serialized voxel sequences. To validate the effectiveness of our approach, we evaluate two model variants: VDM-LION (based on Mamba) and VDM-DSVT (based on Transformer). Experimental results demonstrate that VDM consistently improves performance in both SSMs-based and Transformer-based architectures. Notably, VDM-LION (SSMs) achieves state-of-the-art results on the Waymo, Argoverse 2, and ONCE datasets.

### v3

@@ IV-D Ablation Study
Ablation on Fine-grained Feature Aggregation. To further evaluate the fine-grained feature aggregation capability of VDM, we conducted experiments by refining the voxel resolution. Specifically, we reduced the voxel size along the x and y axes to one-fourth of the original size, while keeping the resolution along the z-axis unchanged, with voxel expansion still enabled.
@@
As reported in Table VI, on the ONCE dataset, the Base model (VDM-OD) exhibits inferior performance across all three categories when only voxel expansion is applied. With the integration of fine-grained feature aggregation (Fine), the model achieves a +1.5 mAP gain, demonstrating the overall effectiveness of finer voxel partitioning.
To delve deeper into these improvements, we analyze the performance per category. It is worth noting that while the simplified VDM-OD (densification-only) variant achieves impressive performance on rigid, large-scale objects such as Vehicles (e.g., 70.5 mAP on nuScenes test set), the full VDM framework demonstrates superior capability in detecting smaller, non-rigid objects. For instance, on the ONCE validation set, the full VDM outperforms VDM-OD by a significant margin on the Pedestrian category (+3.2 mAP in Table V). We attribute this phenomenon to the inherent geometric differences between categories. Large rigid objects (Vehicles) benefit primarily from increased voxel density to recover structural completeness, which VDM-OD provides efficiently. In contrast, small and sparse objects (Pedestrians) possess fewer points and more complex local geometries. The fine-grained aggregation mechanism in the full VDM is essential for capturing these subtle local contexts, which might be overlooked by simple densification. Therefore, although VDM-OD offers a strong baseline for vehicles, the full VDM provides a more balanced and robust solution, particularly for vulnerable road users (VRUs) critical to autonomous driving safety.
V Conclusion
In this paper, we propose the Voxel Densification Module (VDM), a novel pre-serialization stem designed to bridge the spatial-serial gap in existing 3D detectors. By leveraging sparse 3D convolutions, VDM explicitly expands the foreground voxel set to address the sparsity limitation inherent in serialized models, while simultaneously performing fine-grained spatial aggregation to enrich local context. We validate the universality of our approach by integrating VDM into both SSM-based (LION) and Transformer-based (DSVT) architectures. Extensive experiments across four large-scale benchmarks demonstrate that VDM consistently boosts detection performance. Our analysis reveals that while pre-serialization densification serves as the primary driver for mitigating feature sparsity, the fine-grained aggregation mechanism provides specific gains in scenarios requiring richer local geometric details. These results confirm that explicit voxel densification is a generic and effective strategy for advancing point cloud object detection.